### v2

# Tur[k]ingBench: A Challenge Benchmark for Web Agents

Abstract
Recent chatbots have demonstrated impressive ability to understand and communicate in raw-text form. However, there is more to the world than raw text. For example, humans spend long hours of their time on web pages, where text is intertwined with other modalities and tasks are accomplished in the form of various complex interactions. Can state-of-the-art multi-modal models generalize to such complex domains?
To address this question, we introduce TurkingBench, a benchmark of tasks formulated as web pages containing textual instructions with multi-modal context. Unlike existing work which employs artificially synthesized web pages, here we use natural HTML pages that were originally designed for crowdsourcing workers for various annotation purposes. The HTML instructions of each task are also instantiated with various values (obtained from the crowdsourcing tasks) to form new instances of the task. This benchmark contains 32.2K instances distributed across 158 tasks.
Additionally, to facilitate the evaluation on TurkingBench, we develop an evaluation framework that connects the responses of chatbots to modifications on web pages (modifying a text box, checking a radio, etc.). We evaluate the performance of state-of-the-art models, including language-only, vision-only, and layout-only models, and their combinations, on this benchmark. Our findings reveal that these models perform significantly better than random chance, yet considerable room exists for improvement. We hope this benchmark will help facilitate the evaluation and development of web-based agents.111Dataset and baselines are available at: https://turkingbench.github.io
1 Introduction
There has been tremendous progress in building diverse AI models, thanks to the availability of large pre-trained language models (LLM) (Radford et al., 2019). However, the bulk of the progress has been limited to chatbots that process raw text-form in their input or output. However, humans rarely work with text-only interfaces. Instead, they generally like to work with visually situated language, such as web-based applications where language is intertwined with other visual elements such as images and tables. It is only recently that multi-modal chatbots have entered the arena (OpenAI, 2023; Anil et al., 2023; Kil et al., 2024), and plenty is yet to be explored in terms of their capabilities and limits, particularly in the context of web-based exploration. A key limitation is the lack of well-defined and natural benchmarks that encompass the complex, interactive, and multi-modal reasoning for various web-based tasks that provide an objective measure of progress for web-navigating agents.
In this work, we introduce TurkingBench, a benchmark of 158 web-grounded tasks. Examples of these tasks are shown in Figure 1. The information content in these tasks is relayed via various stylistic elements (colors, font sizes, shapes, etc.), structural knowledge (tables, paragraph headings, etc.), as well as modalities (text, images, videos, etc.) to communicate various complex tasks for human users. A thorough understanding of these multi-modal signals is a complex task that requires various skills, including understanding language situated in rich visual cues.
The tasks in TurkingBench are collected from tasks published on the Amazon Mechanical Turk (AMT) crowdsourcing platform, a well-established platform for connecting human workers to solve micro-tasks in return for a reward. We are inspired by the ability humans to solve a variety of tasks through web-based interfaces (Efrat & Levy, 2020).
There are also other distinguishing factors to TurkingBench in comparison to the existing literature. The instructions for solving our tasks are interleaved within each web page, and solving them require effective understanding of the page content. This is distinct from most related benchmark (Deng et al., 2023; Zhou et al., 2023) where the task instructions are provided as as a free-form sentences. Regarding the naturalness of the data, the web-pages in TurkingBench were originally tailored for human crowd workers, making them more natural compared to the recent benchmarks that engineer simplified web pages (Liu et al., 2018; Furuta et al., 2023). Last but not least, since the tasks were originally intended for solving non-trivial annotation tasks on crowdsoucing platforms (e.g., building complex NLP benchmarks) they inherently posses a high degree of complexity. We discuss more related work in §2.
There are a variety of challenges involved in solving our proposed task: (i) Multi-modality—these tasks require the models to “see” or “hear” the task content. (ii) Interaction— The agents’ actions are executed in interaction with the web pages, taking into account the context of the previous changes. (iii) Long context—the web pages (the HTML encoding, screenshots, etc.) are typically longer than the context window of existing models.
We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1). We observe that GPT4 can perform notably higher than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page. However, GPT4’s performance is notably far from the estimated upper bound of the dataset. We also analyze model performance as a function of field type and task length, which identify challenges to our future progress.
In summary, we introduce TurkingBench, a benchmark of web-based tasks that requires multi-modal and interactive reasoning on web pages. We formalize the evaluation framework for interacting with and modifying these pages. The introduction of this benchmark aims to encourage the development of agent models that are not only capable of understanding language within a multi-modal context, but also able to transform their knowledge into a sequence of affordances when interacting with web pages. We evaluate notable models on this benchmark and show that they have noteworthy room for progress. We hope this benchmark motivates and enables further development of assistive, web-based agents.
| MiniWoB++ (Liu et al., 2018) | CompWoB (Furuta et al., 2023) | RUSS (Xu et al., 2021) | WebShop (Yao et al., 2022) | Mind2Web (Deng et al., 2023) | WebArena (Zhou et al., 2023) | TurkingBench (this work) | |
| Inputs | language commands | language commands | customer service queries | shopping queries | language commands | English commands | instructions embedded in web pages |
| The input construction | crowdsourced | manual | mined online | crowdsourced | author-written & LM-augmented | author-written | collected from MTurk |
| Environments | simplified web page | simplified but compositional | real-world websites | shopping web pages | real-world websites | variety of web pages | crowdsourcing websites |
| Natural inputs? | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ |
| Realistic interface? | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ | ✓ |
| Dynamic interaction | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✓ |
| Functional correctness? | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✓ |
2 Related Work
Web-navigation benchmarks.
There are several benchmarks for quantifying abilities of web-navigation agents (Lù et al., 2024; Koh et al., 2024; He et al., 2024). Table Table 1 shows a comparison among a subset of these works. These benchmarks provide an easy way to evaluate models in a reproducible manner. However, all these benchmarks are limited compared to the real-world web. For example, most of the existing datasets operate on websites with limited complexity, such as the ones that are pre-specified, synthesized, or simplified (Liu et al., 2018; Li et al., 2020; Yao et al., 2022; Furuta et al., 2023; Cheng et al., 2023). In almost all the existing benchmarks, the task definitions (instructions) are done after collecting the websites, through manual crowdsourcing or written manually. However, in TurkingBench, instructions are also naturally mined, as they are already embedded in the crowdsourcing web pages. It is also worth noting a limitations of TurkingBench compared to other benchmarks: here the tasks are limited to one page (no page navigation). Given these trade-offs, the field is likely to benefit from all the aforementioned benchmarks.
Instruction-following datasets.
Being able to execute actions on the web also connects to instruction-following literature. Currently, the dominant majority of this literature has has focused on raw text instructions and outputs (Wang et al., 2022; Mishra et al., 2022; Bach et al., 2022; Dwivedi-Yu et al., 2022; Ouyang et al., 2022; Sanh et al., 2022; Shen et al., 2023). The problem of web-exploration (and the proposed benchmark here) is a strict extension of this literature with multiple modalities, with large inputs (long HTML or large screenshots) and strategic decision-making.
Multi-modal reasoning tasks.
TurkingBench is also related to the efforts on multi-modal reasoning. Some works focus on grounding commands to web page elements, setting interaction aside (Pasupat et al., 2018). Several recent efforts study the problem in interactive settings (Gur et al., 2018; Ku et al., 2020; Li et al., 2020; 2022; Li & Li, 2022; Sun et al., 2022; Li et al., 2021; Bai et al., 2021), though their instructions are simple, typically only a few sentences long. Our work differs since our instructions are part of the web pages, and inferring them involves multi-modal reasoning.
Web-based agents.
The idea of intelligent automated assistant agents that collaborates with humans to complete tasks has been around (Allen et al., 2007). One can view web-based agents as extensions of early work on semantic parsing (Das et al., 2010; Clarke et al., 2010; Bordes et al., 2012; Gur et al., 2022). However, general-purpose web-browsing necessitates dealing with challenges well beyond the classic semantic parsing; for example, the diversity of web pages (no two pages are identical), their change over time, the vast action space (a variety of ways that one can interact with web), multi-modality (text, vision, audio), interactive nature of web-browsing, and so on.
The existing literature use a variety of supervision protocols for supervision, for example, behavior cloning (supervised learning of actions) (Gur et al., 2023), reinforcement learning (Liu et al., 2018; Nakano et al., 2021; Humphreys et al., 2022), in-context prompting and demonstrations (Kim et al., 2023; Tao et al., 2023; Sridhar et al., 2023). Here are a few highlighted works, among others: (Sodhi et al., 2023) proposed web agents equipped with hierarchical policies (HeaP), i.e., low-level actions that are guided by high-level plans. (Liu et al., 2023) benchmarks a variety of open-source and open-source and closed-source (API-based) models on several well-known interactive benchmarks. While the focus of our work is introducing a new benchmark, we have intended to provide a diverse variety of models to better understand its difficulty. Our efforts are not exhaustive and leave more room for future work to explore more effective model engineering.
3 TurkingBench: Benchmarking Web Agents with Multi-Modal, Interactive Turking Tasks
We propose the development of TurkingBench. The first purpose of this benchmark is to serve as a standardized evaluation of model generalizability. In addition to supporting evaluation, a subset of this data can streamline model design and development.
3.1 Benchmark Schema
TurkingBench consists of a collection of tasks where each task is a bundle of the following components:
A web template containing instructions for the task, input variables, and input fields for the outputs.
Input values to be instantiated for the input variables.
Annotated output labels provided by crowd workers.
An example is shown is shown in Figure 1. As the example shows, the HTML template contains variables (e.g., ${sys10}) that are instantiated with input values. Furthermore, the HTML template contains input fields for receiving input values from the web page users. For each field, we have values previously collected from crowd workers that we will use for evaluation.
3.2 Collecting the tasks
For a benchmark of tasks grounded in web pages to effectively measure progress and generalizability, it needs to be diverse and with broad coverage. Toward such a recourse, we plan on utilizing two existing sources: (A) We will collect the crowdsourcing templates and the resulting annotations from various AI researchers who are active requesters on AMT. (B) Another incidental resource is the set of tasks published on AMT sandbox,222https://requestersandbox.mturk.com a testing environment for requesters to test out the prototypes of their crowdsourcing templates. Since the sandbox environment contains tasks from all users of AMT, it offers more variety than (A). However, they would require additional work to annotate and clean up.
Ultimately, we put together a collection of 158 crowdsourcing tasks (examples in Figure 1). These include but are not limited to language processing tasks (such as paraphrasing, validating factuality, entailment, sentiments, classification, dialogue quality, and rationale generation) and tasks that involve processing images or videos.
3.3 Statistics
Figure 3 shows the overall statistics of TurkingBench. The dataset contain a large sum of input-output instances (36.2K) across 158 tasks. Furthermore, Figure 2 shows the distribution of the input fields represented in our benchmark. The most common field is “radio,” which is expected as it is commonly used to identify preferences. Here, we also distinguish between “text” input (a small, single-line box) and “textarea” input (a larger, multi-line box for descriptions and paragraphs) fields.
| Measure | Value |
| # of tasks | 158 |
| # of instances | 36.2K |
| avg. # of fields per task | 15.6 |
| avg. length (subwords) of the tasks | 16.8K |
3.4 Programmatic Interaction with Web
We developed a Python library (Figure 4) to streamline the interaction of self-supervised models with our tasks and their evaluations (Figure 4). Our library contains various components. To serve our tasks, we use Turkle333https://github.com/hltcoe/turkle which is an open-source replication of AMT. The content of these tasks are then accessed through the web-browsers that are loaded by Selenium.444https://github.com/SeleniumHQ/selenium
Furthermore, we have developed a library of actions for accessing the pages’ content and making changes to them. This is discussed in §3.4.1. The starting point of interaction is our evaluation script that loops over the tasks and their instances. This is fleshed out in §3.6
3.4.1 A Library of Web Actions
To support model design, we have developed a library of “actions” that can execute various operations on a web page. Our set of actions wrap around the API libraries provided by Selenium (c.f. footnote 4) and PyAutoGUI555https://github.com/asweigart/pyautogui these are sophisticated libraries for web manipulation and the amount of sophistication these libraries provide is significantly more than what is here. Therefore, we build our wrappers around these actions to build an action library with enough complexity to cover the dominant majority of tasks in our benchmark.
| Actions | Modality | Description |
| modify_text | T | modifies the text of input box |
| modify_checkbox | T | modifies the selection of checkbox |
| modify_radio | T | modifies a radio button |
| modify_select | T | selects an item in a drop-down menu |
| modify_range | T | modifies a range input |
| get_html | T | fetches the HTML content of a page |
| capture_screen | V | fetches the screenshot of a page |
| click | V | clicks on a given coordinate |
| type | V | keyboard type at the selected |
| scroll | – | scrolls up or down |
| maximize | – | maximizes the web page |
The list of the actions included in our library is shown in Table 3. In terms of the modalities of information, a subset of these actions are either concerned with executing tasks in text (HTML) modality or visual modality. There are fewer actions in the visual modality since visually many actions are combinations of clicking and typing. This is unlike the text modality where various element types require slightly different treatment.
In terms the nature of their roles, one can split these actions into three groups: (i) Modification actions allow the models to modify each page’s input elements. (ii) Navigation actions allow the models to explore the page, for example, by scrolling up/down or waiting for all elements to be loaded. (iii) Sensing actions allow models to retrieve the latest information about the web page, for example, by fetching its HTML code, taking screenshots of the active page, or around a given element.
Web interaction as “tool” resolution.
The overall task defined here can be viewed as iterative tool resolution based on the context of each task. A web agent solving this task must use a sequence of actions from our action library, based on the context of the task’s instruction (see an example in Figure 5). This connects to the broader literature on Tool-augmented LMs (Schick et al., 2023; Lu et al., 2023; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023) that use Tools to ground LM generations in physical environments. The set of actions defined in our action library is the set of “tools” that can be executed to obtain specific outcomes. The core of this setup is built upon an LM aware of the tools, the problem context, and the ultimate goal.
3.5 Evaluation Metrics
We devise an evaluation metric that is sensitive to the output type:
Text fields: For a given text response, we compute ROUGE metric (Lin, 2004) against each label and compute their maximum value.
Radio/select fields: We compute an exact match of the predicted label against the majority vote of gold labels.
Checkbox fields: The results of the checkbox selection is a set. Therefore, we quantify this as intersection over union metric between the two sets (set of predicted labels and the gold labels).
Range fields: The range predictions provide real or integer values. We compute absolute () distance between the prediction and the labels, averaged across the labels. This is normalized by the largest value among the gold labels to normalize it to range.
The overall score is an average of all the above responses. Given that the field distribution of each type does not vary from one model to another, averaging these different measures of quality is acceptable when comparing different models.
3.6 Evaluation Protocol
Our evaluation protocol consistent of a loop over the various choices of evaluation tasks. Specifically means that, the evaluation script provide the model with the URL where the task can be reached. Furthermore, we provide the model the access to our action library that contains various actions that can be executed by the model.
Furthermore, in this paper we also provide the list of field names on which we will evaluate model responses. This arguable simplifies the task a bit, however, our benchmarks leaves it open for any future analysis with open-ended interaction without any additional hand-holding about the names of the target fields.
Task splits for measuring generalization to unseen instructions.
To accommodate for scenarios that may involve fine-tuning on supervised data, we create different task splits: (i) Train: a set of 121 tasks that can be used for supervised model development. (ii) Test: a set of 20 tasks that be used for evaluating model quality. (iii) Test: a set of 17 tasks that are strictly more difficult than the rest of the tasks, because they require more complex combination of actions or other tasks not covered in Table 3. In this version of the work, we do not provide any experiments on Test since they require more sophisticated engineering that is beyond the scope of this work and should be addressed in future work.
4 Evaluating Models in Solving Web-based Tasks in TurkingBench
We benchmark various models with architectural choices on TurkingBench. We note that, our goal here is to provide reasonable baselines for our proposed benchmark and hence, we avoid specialized models that may use specific assumptions or supervision of our data.
Models.
As discussed earlier, our models need to consume information on a mix of information modalities. A model can, therefore, consume the instructions as text, image, or a combination of both. We experiment with state-of-the-art models like GPT-4 and Claude-2.1.666Accessed via their APIs in January 2024. For GPT-4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information. When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page. (2) “relevant” indicates a few lines of HTML above and under the input field the model is currently solving. To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation task. This adds up to web pages evaluated with a total of roughly input fields evaluated.
We note that, besides these general-purpose models there are more specialized text-based models that are specialized for web-exploration either through pre-training on text data (Aghajanyan et al., 2021; 2022; Gur et al., 2022), specialized models for analyzing web-based components (Huang et al., 2022; Tao et al., 2022; Ebrahimi et al., 2022; Chen et al., 2022) or visual perception of web (Dosovitskiy et al., 2020; Rust et al., 2022; Lee et al., 2022). We delegate such explorations that are likely to yield better results to the future.
An oracle model for ceiling performance.
We implement an oracle baseline that mimics the gold labels for each input. This data provides granular action sequences that facilitate future model developments. These action sequences are extracted by connecting crowdworker annotations to the name of the corresponding HTML input elements in task web pages and executing an appropriate action similar to the ones shown in Figure 5. A pseudo-code of the oracle model is shown in Alg.2.777See the a screencast of the oracle model’s execution. We note that, the oracle baseline is limited to tasks that do not require complex annotations (such as drag-and-drop) that were included in Test (discussed in our evaluation split §3.5).
A “do-nothing” model as a lower-bound.
We also evaluate a trivial baseline that executes nothing (hence the name, “do-nothing”). As we see later in the results, there are tasks that score more that zero. This is because there are tasks that ask, for example, corrections on a given text where at times doing nothing is the right action. Despite such settings, we will see the overall there is a big margin between the do-nothing baseline and the oracle baseline.
4.1 Empirical Results
We outline our empirical observations on our proposed benchmark.
Modalities |
Score (%) | |||||||
| Do-nothing | – | – | – | – | 7.8 | |||||||
| GPT4 | T | relevant | 7 | 5k | 21.3 | |||||||
| T | full | 7 | 82k | 39.3 | ||||||||
| GPT4-V | T+ V | relevant | 7 | 8.4k | 24.5 | |||||||
| T+ V | full | 7 | 86k | 41.7 | ||||||||
| Claude2.1 | T | full | 7 | 82k | 22.6 | |||||||
| Oracle | – | – | – | – | 100.0 |
Despite the remarkable performance of generalist models, they remain far from our ceiling performance.
We show the results of evaluating our main models in Table 4. For each row, we indicate whether the input to the model contains text-only (T) vs. text-vision (T+V). All the models show have 7 in-context demonstrations of tasks and desired action. Note in the table we also show the size input prompts measured in subwords. The best performance is obtained by GPT-4 vision-language model (T+V) with access to full text of each task. This configuration also happens to have a extremely large prompt length (86k subwords) and it shows the remarkable ability of this model to exploit long-range dependencies. We note that the gains of the vision model (T+V) over text-only model (T) is minimal ( vs. ).
Text-only and vision-language models show complementary capability.
For our top configurations in Table 4 (GPT-4, 7 demonstrations with “full” HTML encoding) we show the breakdown of the model performances across various tasks in Figure 6. On 12 evaluation tasks (out of 20) the performance of these two models are notably different from one another. While it is intuitive to speculate that task with richer task interface likely benefit more from visual, in practice it was not clear what exact details details about task design leads to such differences.
The gains of in-context demonstrations plateau quickly.
As mentioned, we use few-shot prompting of models in order to steer them their predictions. Here we study the effect of the number of demonstrations in model performance. As the results are shown in Figure 7, the gains of in-context demonstrations quickly plateaus when the number of demonstrations just above 3 demonstrations.
Performance across field type.
To better understand the difficulty of the task as a function of field types, we show the performance of GPT-4 (text input; 7 demonstrations) per input field type in Figure 8. We distinguish between “text” and “text area” because they’re defined with different tags in HTML (<input type="text"> vs. <textarea>). The former is often for short inputs and the latter for longer, multi-sentence outputs. This is also refelcted in lower performance results for “textarea” inputs.
5 Conclusion
We introduced TurkingBench to facilitate research on general-purpose web-based agents. Our benchmark is focused on tasks defined in the context of web pages, such as those that tend to appear on crowdsourcing platforms. Our benchmark includes a comprehensive Python-based framework that facilitates evaluation and model development. We hope this benchmark facilitates further development of web-based assistant agents.
Limitations.
There are various limitations in our work. (i) Despite the complexity of crowdsourcing tasks, they make up just a tiny fraction of the web. We are willing to make this trade-off to obtain natural tasks with ground-truth labels. Other existing benchmarks often need to develop artificial (manually-annotated) tasks and manually annotate gold labels for them. (ii) Because of the choice of domain, our tasks do not involve interactions that require jumping over multiple web pages—instead, all the tasks are contained within a given web page. Despite these simplifications, as we see in our experiments, the task remains challenging for the models. (iii) Our evaluation setup provides some hand-holding for the models by providing the names of the fields that need to be modified. Future work should study variants of our task where such hints are not provided to the model.
Author Contributions
KX led the major implementation of the evaluation framework and obtaining the final results. YK implemented the first version of the repository. KS, YW, AB participated in meetings, provided feedback and helped with the draft. JZ helped with the paper draft. DK lead the project, implementation and writing.
Acknowledgements
This project is supported by a generous gift from Amazon. The authors would like to thank Chris Callison-Burch, Mark Dredze, Karthik Narasimhan, Nikhil Sharma, Owen Bianchi, and Elizabeth Salesky for their constructive discussions. GPU machines for conducting experiments are provided by the ARCH (Rockfish) cluster (https://www.arch.jhu.edu).
6 Ethics Statement
A notable portion of humans spend more time on the web than the outside world (Perrin & Kumar, 2019). Consider an academic individual who spend long hours of their day responding to emails, writing proposals, scheduling talks, setting up meetings, etc. As a result, AI-driven automation of web navigation has the potential to simplify the experience of many and make web more accessible (Deng, 2023).
We acknowledge a concern that this work may enable technologies that would replace crowd workers, who have been an essential part of the AI development lifecycle. The concern is not specific to this work but applies broadly to all AI developments. We note that all the models discussed here are mainly enabled by developments outside this benchmark. Furthermore, the experimental results on our benchmark show that state-of-the-art self-supervised models are still far from the benchmark’s ceiling performance, even on our (simplified) evaluation setup. This means it will take much more progress before AI systems can fully solve autonomous crowdsourcing.
We hope this work enables benevolent use-cases of AI, such as AI systems that help improve the quality and productivity of crowd workers.